# Robust and Computation-Aware Gaussian Processes

### v1

@@ F.1 CaGP Lacks Robustness
Based on [2], the PIF leads to the following form:
@@ F.2 Mean convergence of RCaGP
We first show the corresponding empirical-risk minimization problem of RCGP whenever ${\bf{m}}=\mathbf{0}$. Following [2, proof of Proposition 3.1], we can rewrite $L_{n}^{w}$ and formulate the RCGP objective as follows:
@@
The optimization problem in S119 allows us to apply the representer theorem [36]. It implies that the solution of S119 can be written as a weighted sum, i.e.,
@@
Since the objective in S123 is a convex function of ${\boldsymbol{\alpha}}$, we find that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the minimum of the objective (S119 and S123). Furthermore, we can verify that $L_{n}^{w}$ is a convex function w.r.t. ${\bf{f}}$. Therefore, we conclude that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the unique solution to S119. As a remark, Proposition F.1 closely connects with [20, Theorem 3.4].
@@
Relative bound errors. We first provide the equivalence of [45, Proposition 2]:
@@
The proof is by directly substituting ${\bf{C}}_{i},\hat{{\bf{K}}},{\bf{v}}_{\ast}$ in [45] with $\tilde{{\bf{C}}}_{i},\tilde{{\bf{K}}},\hat{{\bf{v}}}$, respectively. Next, we establish the convergence theorem of RCaGP through the following proposition:
@@
The third inequality stems from the definition of ${\bf{J}}_{\bf{w}}$ and the fact that the maximum eigenvalue of a diagonal matrix is the largest component of its diagonal. Assuming $\hat{\mu}_{i}(.)=m(.)+\sum_{j=1}^{n}(\tilde{{\bf{v}}}_{i})_{j}k(.,{\bf{x}}_{j})=m(.)+k(.,{\bf{X}})\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and applying result from [45], we have that
@@ Appendix G Details of Expert-guided Robust Mean Prior
[…]
@@
$\qquad\quad\;\;=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})\left(\bar{\delta}_{o}(\bar{o}_{o})+(1-\bar{\delta}_{o})(1-\bar{o}_{o})\right)$
$\qquad\quad\;\;=\mathcal{B}(\alpha=\alpha_{o}+\bar{o}_{o},\beta=\beta_{o}+1-\bar{o}_{o}),$ (S136)
@@
[…]
@@
$\qquad\quad\;\;={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})\,{\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ (S141)
$\qquad\quad\;\;={\mathcal{N}}\left(\mu=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},\sigma^{2}=(\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}})^{-1}\right),$ (S142)
@@
which equals the posterior mean. The algorithm to implement the informative mean prior is presented in Algorithm 1.
@@ H.1 UCI regression datasets
The dataset consists of $n=506$ observations, each representing a suburban or town area in Boston. It encompasses $d=13$ features containing data like the average number of rooms in dwellings, pupil-teacher ratios, and per capita crime rates. We try to predict the median price of homes residents own (excluding rented properties). The dataset can be found at https://www.cs.toronto.edu/~delve/data/boston/bostonDetail.html.
@@ H.2 High-throughput Bayesian Optimization problems
The widely used Hartmann benchmark function [38].
@@
The goal of this task is to find an optimal $12$-dimensional control policy that allows an autonomous lunar lander to consistently land without crashing. The final objective value we optimize is the reward obtained by the policy averaged over a set of 50 random landing terrains. For this task, we use the same controller setup used by [8].
@@
The rover trajectory optimization task introduced by [44] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a set of obstacles. We use the same obstacle set up as in [29].
@@
We optimize the $180-$dimensional DNA task from the LassoBench library [49] of benchmarks based on weighted LASSO regression.
@@ H.3 Additional outlier contamination protocols on UCI regression datasets
Our outlier contamination protocols follow the same settings as [2]. We recall the asymmetric outlier case described in the main text in Section 7.

### v2

@@ F.1 CaGP Lacks Robustness
Based on [altamirano2023robust], the PIF leads to the following form:
@@ F.2 Mean convergence of RCaGP
We first show the corresponding empirical-risk minimization problem of RCGP whenever ${\bf{m}}=\mathbf{0}$. Following [altamirano2023robust, proof of Proposition 3.1], we can rewrite $L_{n}^{w}$ and formulate the RCGP objective as follows:
@@
The optimization problem in S119 allows us to apply the representer theorem [scholkopf2001generalized]. It implies that the solution of S119 can be written as a weighted sum, i.e.,
@@
Since the objective in S123 is a convex function of $\boldsymbol{\alpha}$}, we find that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the minimum of the objective (S119 and S123). Furthermore, we can verify that $L_{n}^{w}$ is a convex function w.r.t. $\bf{f}$}. Therefore, we conclude that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the unique solution to S119. As a remark, Proposition F.1 closely connects with [kanagawa2018gaussian, Theorem 3.4].
@@
Relative bound errors. We first provide the equivalence of [wenger2022posterior, Proposition 2]:
@@
The proof is by directly substituting ${\bf{C}}_{i},\hat{{\bf{K}}},{\bf{v}}_{\ast}$ in [wenger2022posterior] with $\tilde{{\bf{C}}}_{i},\tilde{{\bf{K}}},\hat{{\bf{v}}}$, respectively. Next, we establish the convergence theorem of RCaGP through the following proposition:
@@
The third inequality stems from the definition of ${\bf{J}}_{\bf{w}}$ and the fact that the maximum eigenvalue of a diagonal matrix is the largest component of its diagonal. Assuming $\hat{\mu}_{i}(.)=m(.)+\sum_{j=1}^{n}(\tilde{{\bf{v}}}_{i})_{j}k(.,{\bf{x}}_{j})=m(.)+k(.,{\bf{X}})\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and applying result from [wenger2022posterior], we have that
@@
F.3 SVGP with Relevance Pursuit
Proposition F.4.
Let ${\mathcal{D}}_{\setminus i}=\{({\bf{x}}_{j},y_{j}):j\neq i\},{\bf{Z}}={z_{m}}$ is the inducing points, ${\boldsymbol{\rho}}={\boldsymbol{\rho}}_{\setminus i}+\rho_{i}\mathbf{e}_{i}$, where ${\boldsymbol{\rho}},{\boldsymbol{\rho}}_{\setminus i}\in_{+}^{n}$, ${\boldsymbol{\rho}}_{\setminus i}]_{i}=0$, and $\mathbf{e}_{i}$ is the ith canonical basis vector. Then keeping ${\boldsymbol{\rho}}_{\setminus i}$ fixed,
$\rho_{i}^{\ast}={\mathrm{argmax}}_{\rho_{i}}\,\mathrm{ELBO}({\boldsymbol{\rho}}_{\setminus i}+\rho_{i}\mathbf{e}_{i})=\left[\frac{1}{\sqrt{{\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}}}(y_{i}-{\mathbb{E}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}])-{\mathbb{V}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}]\right]_{+}$ (S134)
where $y_{i}=f({\bf{x}}_{i})+\epsilon_{i}$. The quantities can be expressed as a function of ${\bf{A}}^{-1}=({\bf{Q}}_{\bf{f}}+{\bf{D}}_{{\sigma_{\mathrm{noise}}^{2}}+{\boldsymbol{\rho}}})^{-1}$ :
${\mathbb{E}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}]=y_{i}-[{\bf{A}}^{-1}{\bf{y}}]_{i}/[{\bf{A}}^{-1}]_{ii}\quad\text{and}\quad{\mathbb{V}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}]=1/[{\bf{A}}^{-1}]_{ii}$ (S135)
where ${\bf{D}}_{{\sigma_{\mathrm{noise}}^{2}}+{\boldsymbol{\rho}}}$ is a diagonal matrix whose entries are ${\sigma_{\mathrm{noise}}^{2}}+{\boldsymbol{\rho}}$.
Proof: Following [titsias2009variational], the evidence lower bound (ELBO) of GP can be written as
$2\mathrm{ELBO}({\boldsymbol{\theta}})=n\log(2\pi)+\log\det(\mathbf{A})+{\bf{y}}^{\top}{\bf{A}}^{-1}{\bf{y}}+{\mathrm{tr}}({\bf{D}}_{{\boldsymbol{\rho}}+{\sigma_{\mathrm{noise}}^{2}}}^{-1}({\bf{K}}-{\bf{Q}}_{\bf{f}})),$ (S136)
where ${\bf{A}}={\bf{Q}}_{\bf{f}}+{\bf{D}}_{{\boldsymbol{\rho}}+{\sigma_{\mathrm{noise}}^{2}}}$, and ${\bf{Q}}_{\bf{f}}={\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}$
Following [ament2024robust], we partition matrix $\bf{A}$} to separate the effect of $\rho_{i}$ and use Schur’s complement:
${\bf{A}}^{-1}=\begin{bmatrix}{\bf{A}}^{-1}_{\setminus i}+\hat{{\bf{u}}}\beta_{i}\hat{{\bf{u}}}^{\top}&-\hat{{\bf{u}}}\beta_{i}\\ -\hat{{\bf{u}}}^{\top}\beta_{i}&\beta_{i}\end{bmatrix}$, (S137)
where
${\bf{A}}_{\setminus i}=[{\bf{Q}}_{\bf{f}}]_{\setminus i}+{\bf{D}}_{{\boldsymbol{\rho}}_{\setminus i}+{\sigma_{\mathrm{noise}}^{2}}},$ (S138)
$\hat{{\bf{u}}}={\bf{A}}^{-1}_{\setminus i}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}),$ (S139)
$\hat{\beta}_{i}=([k({\bf{x}}_{i},{\bf{x}}_{i})+{\sigma_{\mathrm{noise}}^{2}}+\rho_{i}]-k({\bf{x}}_{i},{\bf{Z}}){\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}{\bf{A}}_{\setminus i}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}))^{-1}.$ (S140)
Quadratic term:
${\bf{y}}^{\top}({\bf{A}}+{\bf{D}}_{{\boldsymbol{\rho}}+{\sigma_{\mathrm{noise}}^{2}}})^{-1}{\bf{y}}={\bf{y}}^{\top}_{\setminus i}{\bf{A}}^{-1}_{\setminus i}{\bf{y}}_{\setminus i}+\beta_{i}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}.$ (S141)
Determinant term:
$\det({\bf{A}})=\det({\bf{Q}}_{\bf{f}}+{\bf{D}}_{{\boldsymbol{\rho}}_{i}+{\sigma_{\mathrm{noise}}^{2}}})=\det({\bf{A}}_{\setminus i})\beta_{i}^{-1}.$ (S142)
Following [ament2024robust], we formulate the ELBO difference as
$2(\mathrm{ELBO}({\boldsymbol{\rho}})-\mathrm{ELBO}({\boldsymbol{\rho}}_{\setminus i}))=-\hat{\beta}_{i}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}-\log(\hat{\beta}_{i}^{-1})-\rho_{i}({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}).$ (S143)
The derivative of the difference in ELBO w.r.t. $\rho_{i}$, is
$\partial_{\rho_{i}}2(\mathrm{ELBO}({\boldsymbol{\rho}})-\mathrm{ELBO}({\boldsymbol{\rho}}_{\setminus i}))=({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}\hat{\beta}_{i}^{2}-\hat{\beta}_{i}-({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}).$ (S144)
Using the quadratic formula, we obtain the solution $\hat{\beta}_{1,2}=\frac{1\pm\sqrt{1+4({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})}}{2({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}}$. Note that $({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}>0$ and ${\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}\geq 0$. Therefore, the smaller root $\rho_{2}$ is negative when $({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})>0$ and zero when $({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})=0$. Based on the $\hat{\beta}$ formulation (predictive variance cannot be zero), we reject $\hat{\beta}_{2}<0$. Similarly, we reject $\hat{\beta}_{2}=0$, following [ament2024robust]. By ignoring the constant term in $\hat{\beta}_{2}$ and following the approach of [ament2024robust], we solve $\hat{\beta}_{2}^{-1}=1/\sqrt{({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})$ for $\rho_{i}$ and projecting to non-negative half-line, we get
$\rho_{i}=[\frac{1}{\sqrt{({\bf{K}}_{ii}}-[{\bf{Q}}_{\bf{f}}]_{ii})}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})-(k({\bf{x}}_{i},{\bf{x}}_{i})+{\sigma_{\mathrm{noise}}^{2}}-k({\bf{x}}_{i},{\bf{Z}}){\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}{\bf{A}}_{\setminus i}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}))]_{+}.$ (S145)
Lastly, we note that
${\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i}={\bf{y}}_{\setminus i}^{\top}({\bf{A}}^{-1}_{\setminus i}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}))-y_{i}={\mathbb{E}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}]-y_{i},$ (S146)
$k({\bf{x}}_{i},{\bf{x}}_{i})+{\sigma_{\mathrm{noise}}^{2}}-k({\bf{x}}_{i},{\bf{Z}}){\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}{\bf{A}}_{\setminus i}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}})={\mathbb{V}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}].$ (S147)
Following [Rasmussen2006], we have that
${\mathbb{E}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}]=y_{i}-[{\bf{A}}^{-1}{\bf{y}}]_{i}/[{\bf{A}}^{-1}]_{i},$ (S148)
${\mathbb{V}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}]=1/[{\bf{A}}^{-1}]_{ii},$ (S149)
which follow the LOO predictive values.
@@ Appendix G Details of Expert-guided Robust Mean Prior
$p(\bar{\delta}_{o})=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})$ $\text{prior},$ (S150)
$p(\bar{o}_{o}\,|\,\bar{\delta}_{o})=\bar{\delta}_{o}\bar{o}_{o}+(1-\bar{\delta}_{o})(1-\bar{o}_{o})$ $\text{likelihood},$ (S151)
@@
$\hskip 17.00024pt\qquad\;\;=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})\left(\bar{\delta}_{o}(\bar{o}_{o})+(1-\bar{\delta}_{o})(1-\bar{o}_{o})\right)$
$\hskip 17.00024pt\qquad\;\;=\mathcal{B}(\alpha=\alpha_{o}+\bar{o}_{o},\beta=\beta_{o}+1-\bar{o}_{o}),$ (S152)
@@
$p(\bar{\mu}_{o})={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})$ $\text{prior},$ (S154)
$p(\bar{y}_{o}|\bar{\mu}_{o})={\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ $\text{likelihood},$ (S155)
@@
$\hskip 17.00024pt\qquad\;\;={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})\,{\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ (S157)
$\hskip 17.00024pt\qquad\;\;={\mathcal{N}}\left(\mu=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},\sigma^{2}=(\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}})^{-1}\right),$ (S158)
@@
which equals the posterior mean. The algorithm to implement the informative mean prior is presented in Algorithm˜1.
@@ H.1 UCI regression datasets
The dataset consists of $n=506$ observations, each representing a suburban or town area in Boston. It encompasses $d=13$ features containing data like the average number of rooms in dwellings, pupil-teacher ratios, and per capita crime rates. We try to predict the median price of homes residents own (excluding rented properties). The dataset can be found at https://www.cs.toronto.edu/˜delve/data/boston/bostonDetail.html.
@@ H.2 High-throughput Bayesian Optimization problems
The widely used Hartmann benchmark function [surjanovic2020virtual].
@@
The goal of this task is to find an optimal $12$-dimensional control policy that allows an autonomous lunar lander to consistently land without crashing. The final objective value we optimize is the reward obtained by the policy averaged over a set of 50 random landing terrains. For this task, we use the same controller setup used by [eriksson2019scalable].
@@
The rover trajectory optimization task introduced by [wang2018batched] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a set of obstacles. We use the same obstacle set up as in [robot].
@@
We optimize the $180$-dimensional DNA task from the LassoBench library [sehic2022lassobench] of benchmarks based on weighted LASSO regression.
@@ H.3 Additional outlier contamination protocols on UCI regression datasets
Our outlier contamination protocols follow the same settings as [altamirano2023robust]. We recall the asymmetric outlier case described in the main text in Section 7.